Tuesday, July 10, 2007
Dental School :
- Ceiling grid installation is nearly complete on the 7th and 8th floors of the Clinic Building and continues on the 3rd and 4th floors of the Administration Building.
- Sprinkler head adjustments continue to be made on the 7th and 8th floors of the Clinic Building.
- Sheetrock installation around the 5th floor linear diffusers is nearly complete.
- Viable sample results on air handling units E8A, E8C, and EL3B failed.
- Zimmer is re-cleaning these units and re-samples are to be taken this evening.
- Units E8B and E8D passed and were turned on this afternoon.
- The new network cables continue to be terminated throughout the Clinic Building floors.
- The cable trays on part of the 2nd floor of the Administration Building were raised to accommodate the drop ceiling.
- Carpet installation on the 2nd, 3rd and 4th floors of the Clinic Building continued today.
- Both of the sections of the new cooling tower were lifted onto the roof of the Power Plant and set on the prepared steel structure.
1542 Tulane Ave. :
- Ordes pulled wiring from panel PLL-8 to the disconnect box that will supply power to the guard station once the wiring is completed.
- Arc Mechanical is fabricating and installing new pipe on the discharge side of the condensate receiver.
- Stratos Elevator received the new elevator cab for elevator #2 and is dismantling the existing cab.
- The contractor is installing new chain link fencing in Student Lot #3 along S. Prieur Street.
- All of the concrete wheel-stops have been installed and the striping is completed in Lot #3.
- The contractor is applying the 1st layer of asphalt in Lot #2.
- The contractor is also setting the wood forms for the new sidewalk adjacent to Lot #2.
- The installation of sod around the CSRB and the Lions/LSU Clinic is complete.
- The installation of plants in the large plant bed on the Gravier St. side of the Lions/LSU Clinic is progressing.